A pump is used to pump water at 60oF from a reservoir whose surface is 4.0...

  1. A pump is used to pump water at 60oF from a reservoir whose surface is 4.0 ft above the centerline of the pump inlet.  The piping system from the reservoir to the pump consists of 10.5 ft of cast iron pipe with a diameter of 4.0 in and an average roughness (ε) of 0.02 inches.  There are several minor losses:  a sharp edged inlet (K – 0.5), three flanged smooth 90o regular elbows (K – 0.3 each), and a fully open flanged globe valve (K – 6.0).  

Estimate the maximum volume flowrate (Q) that can be pumped without cavitation if the NPSHRequired is 6.5 ft  (assume Pv = 0.46 psia; Patm = 14.7 psia).  

Assume fully turbulent flow!

Answer (Q = 596.95 gpm)
Answer #1

Ans) We know,

zmax = (Patm - Pv) / - Hf - Hm - NPSHr .......................(1)

where, zmax = Maximum elevation of pump to avoid cavitation = 4 ft

Patm = Atmospheric pressure = 14.7 psi or 2116.8 psf

Pv = Vapor pressure of water = 0.46 psi or 66.24 psf

Hf = Frictional head loss

Hm = Minor head loss

Friction loss, Hf = f L V2 / 2 g D

For fully developed turbulent flow , f = 1/ [1.14 + 2 Log (D/e)]2

where, e is roughness of pipe = 0.02 in

=> f = 1 / [1.14 + 2 Log(4/0.02)]2

=> f = 0.030

=> Hf = 0.030(10.5) V2 / (2 x 32.2 x 0.333)

=> Hf = 0.014 V2

Minor loss, Hm = K V2 / 2 g

=>[0.50 + 3(0.30) + 6] V2 / 2 g

=> Hm = 7.4 V2 / (2 x 32.2)

   => Hm = 0.114 V2

Putting values in equation 1,

=> 4 = (2116.8 - 66.24)/62.4 - 0.014 V2 - 0.114 V2 - 6.5

=> 4 = 32.86 - 0.12958 V2 - 6.5

=> 0.128 V2 = 22.36

=> V = 13.21 ft/s

We know, flow rate (Q) = A x V

Area of pipe,A = (/4) (0.333)2 = 0.08705 ft2

=> Q = 0.08705 x 13.21 = 1.15 cfs 596 gpm